Rice A is being sold at 22 dollars per kilo while rice B is being sold at 27 dollars per kilo. How many kilograms of rice A should be mixed with rice B to make a new sack of rice that can be at 25 dollars per kilo.









Answer by ankor@dixie-net.com(22740)

Rice A is being sold at 22 dollars per kilo while rice B is being sold at 27 dollars per kilo. How many kilograms of rice A should be mixed with rice B to make a new sack of rice that can be at 25 dollars per kilo.
:
We need to know how many kilos is the new sack of rice, to do this.
:
Let's assume the resulting $25 sack of rice weighs 100 kilos
let x = amt of $27 rice
then
(100-x) = amt of $22 rice
:
A typical mixture problem
27x + 22(100-x) = 25(100)
27x + 2200 - 22x = 2500
27x - 22x = 2500 - 2200
5x = 300
x = 300/5
x = 60 kilos of $27 rice (B)
and
100 - 60 = 40 kilos of $22 rice (A)